OpenStudy (anonymous):

the areas of two similar rectangles are 192 cm^2 and 363 cm^2. what is the similarity ratio of the two rectangles? need an answer ASAP!!!!

OpenStudy (anonymous):

the ratio of the areas will be: \[\frac{192}{363} = \frac{64}{121}\] The ratio of the sides will be the square root of that: \[\sqrt(\frac{64}{121})= \frac{8}{11}\]

OpenStudy (anonymous):

i will stick to the similarities of the ares=64/121 or 121/64

OpenStudy (anonymous):

areas i mean..lol

OpenStudy (anonymous):

lol, it seemed kinda "convenient" that they happen to reduce to perfect squares haha

OpenStudy (anonymous):

can you answer the rest of my questions ?